# Rings of non-commutative functions and their fields of fractions

### v1

@@ Abstract
We show, in particular, that the ring of (uniformly) entire non-commutative (NC) functions in $d\in\mathbb{N}$ non-commuting variables, $\mathscr{O}_{d}$, is a semifir. Every finitely–generated right (or left) ideal in $\mathscr{O}_{d}$ is closed, which yields an analytic extension of G. Bergman’s nullstellensatz for the free algebra. Any semifir admits a universal skew field of fractions; applying this to $\mathscr{O}_{d}$ yields the universal skew field of “NC meromorphic expressions”, $\mathscr{M}_{d}$. We show that any $f\in\mathscr{M}_{d}$ has a well-defined domain and evaluations in a large class of stably-finite topological algebras, including finite $C^{*}$-algebras, extending a result of Cohn for NC rational functions. As an application, we extend the almost sure convergence result of Haagerup and Thorbjørnsen for free polynomials evaluated on tuples of random matrices to the setting of NC meromorphic expressions.
@@ 1 Introduction
Non-commutative function theory originated in the work of J.L. Taylor in multivariate spectral theory and functional calculus for tuples of non-commuting operators [48, 47]. A non-commutative function in several non-commuting variables is a function on tuples of square matrices of any fixed, finite size, taking values in square matrices, and which has natural invariance properties enjoyed by free polynomials, $p\in\mathbb{C}\langle\mathfrak{z}\rangle=\mathbb{C}\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle$. Here, $\mathbb{C}\langle\mathfrak{z}\rangle$ denotes the free algebra over $\mathbb{C}$ on $d$ generators. We will call elements of $\mathbb{C}\langle\mathfrak{z}\rangle$ non-commutative or free polynomials. A function, $f$, on a direct sum-closed subset of the complex NC universe of $d$-tuples of square matrices over $\mathbb{C}$ of any fixed finite size is an NC function if it: (i) respects the grading (matrix size), (ii) respects direct sums and (iii) respects joint similarities, see Subsection 2.1 for definitions and details. It is readily checked that any $p\in\mathbb{C}\langle\mathfrak{z}\rangle$ is an NC function on the entire NC universe. These three axioms defining NC functions are surprisingly rigid: Any NC function that is ‘continuous’ or even ‘locally bounded’ in an ‘open’ NC set is automatically holomorphic in the sense that it is Fréchet differentiable at any point in its domain and analytic in the sense that it has a convergent Taylor-type power series expansion about any point of its domain with non zero ‘radius of convergence’. (In order to precisely define the punctuated terms in the previous sentence, we need to introduce a topology on the NC universe, which we will do in Subsection 2.1. The topology that is most relevant for our investigations here is called the uniform topology [25, Section 7.2].)
@@
where $\omega=i_{1}\cdots i_{n}$ is any word comprised of letters $i_{j}$, chosen from the alphabet, $i_{j}\in\{1,\cdots,d\}$. Here, $\mathbb{W}_{d}$ denotes the free monoid, the set of all such words in $d$ letters. We also consider the empty word, $\emptyset$, consisting of no letters, and the free monomials, $\mathfrak{z}^{\omega}$, are then defined in the obvious way: $\mathfrak{z}^{\omega}=\mathfrak{z}_{i_{1}}\cdots\mathfrak{z}_{i_{n}}$ if $\omega=i_{1}\cdots i_{n}$ and $\mathfrak{z}^{\emptyset}=:1$. G. Popescu [41] introduced an analogue of the Cauchy–Hadamard radius of convergence formula for such free FPS, $f$, given by the formula
@@
see also [29] and [25]. In the above, $|\omega|\in\mathbb{N}\cup\{0\}$ denotes the length of $\omega\in\mathbb{W}_{d}$, i.e., the number of letters it contains. Given any $R\in(0,+\infty]$, we study the rings, $\mathscr{O}_{d}(R)$, of all free formal power series with radius of convergence at least $R$. We write $\mathscr{O}_{d}:=\mathscr{O}_{d}(+\infty)=\cap_{R>0}\mathscr{O}_{d}(R)$, for the ring of NC entire functions. As proven in [41, Theorem 1.1], if $f\in\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$, has radius of convergence $R_{f}>0$, then $f(X)$ converges absolutely and uniformly in an NC row-ball, $R_{f}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, centred at the origin, $0=(0,\cdots,0)\in\mathbb{C}^{1\times d}$, of the $d$-dimensional NC universe, consisting of all square matrix $d$-tuples, $X=(X_{1},\cdots,X_{d})\in\mathbb{C}^{n\times n}\otimes\mathbb{C}^{1\times d}$, which define bounded linear maps from $\mathbb{C}^{n}\otimes\mathbb{C}^{d}$ into $\mathbb{C}^{n}$, with norm less than $R_{f}$. Hence, any $f\in\mathscr{O}_{d}(R)$ defines a (uniformly) analytic NC function in $R_{f}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. In particular, $\mathscr{O}_{d}$ consists of holomorphic and analytic non-commutative functions on the entire NC universe and is, in a very natural sense, a canonical multivariate extension of the ring of entire functions in $\mathbb{C}$ to several non-commuting variables.
In pure non-commutative algebra, semi-free ideal rings were introduced by P.M. Cohn in his study of localizations and universal skew fields of fractions of non-commutative rings, see [10] for a full development of this theory. A ring, $\mathscr{R}$, is a semi-free ideal ring (semifir for short), if every finitely generated right ideal of $\mathscr{R}$ is free as a right $\mathscr{R}$-module. (This notion is left-right symmetric.) Cohn showed that every semi-free ideal ring admits a (necessarily unique) universal skew field of fractions. There are not many known examples of semi-free ideal rings, and many non-commutative rings do not have universal localizations. (Certainly any ring with zero divisors cannot be embedded into a skew field.) Among the few known examples are the free algebra, the ring $\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle=\mathbb{C}\langle\!\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle\!\rangle$ of formal power series in several non-commuting variables, the sub-ring of algebraic formal power series, the group ring of the free group and group rings over locally free groups (see, for example, [10] and [14]). Some more recent examples of universal localizations and of semifirs were constructed by Klep, Vinnikov, and Volčič in [26, 27]. The first main result of this paper provides a new family of examples of semifirs of uniformly analytic NC functions:
@@
It follows that each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$ has a universal skew field of fractions, which we denote by $\mathscr{M}_{d}(R)$. In particular, we will call $\mathscr{M}_{d}:=\mathscr{M}_{d}(\infty)$ the field of (global) NC (uniformly) meromorphic expressions – c.f. the local field of (uniformly) meromorphic NC germs at $0\in\mathbb{C}^{1\times d}$ investigated in [26]. Our proof of the above theorem relies heavily on the operator–algebraic “Non-commutative Hardy space theory” initiated by Popescu and Davidson–Pitts, in which the full Fock space over $\mathbb{C}^{d}$ is interpreted as the “free Hardy space” of square–summable formal power series in $d$ NC variables [41, 40, 38, 13]. In particular, we will employ the non-commutative inner-outer factorization [13, 34, 36] for the free Hardy space, and extensions of methods in [24] in conjunction with the structure of $\mathscr{O}_{d}(R)$ as an inverse limit of non-commutative disk algebras on balls of increasing radii [5].
We further obtain an analytic extension of a non-commutative Nullstellensatz for the free algebra due to G. Bergman to each of our semifirs, $\mathscr{O}_{d}(R)$, see [23]. To this end, define the directional variety of $S\subseteq\mathscr{O}_{d}(R)$ as
@@
In Section 6, we extend a theorem of P.M. Cohn for evaluation of non-commutative rational functions in any stably-finite algebra. Here, the free skew field of all NC rational functions, $\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,:=\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$, is the universal skew field of fractions of the free algebra, $\mathbb{C}\langle\mathfrak{z}\rangle$ [4, 10]. Namely, we prove that NC meromorphic elements in $\mathscr{M}_{d}$ have well-defined evaluations and domains in any stably finite algebra, $\mathscr{A}$, that admits well-defined unital ring homomorphisms $\varphi:\mathscr{A}\rightarrow\mathscr{O}_{d}$; in particular in any stably finite $C^{*}$-algebra. This result enables us to develop concrete applications of our results to D.-V. Voiculescu’s free probability theory and random matrix theory. Namely, we extend the results of Haagerup and Thorbjørnsen for the free algebra [20] and of Yin [54] for the free skew field, to the larger field of NC meromorphic expressions. To describe these results in more detail, let $s_{1},\cdots,s_{d}$ denote the real parts of the left creation operators on the full Fock space over $\mathbb{C}^{d}$. This $d$-tuple is the prototypical example of free semicircular random variables in free probability theory.
@@
Let $(\Omega,P)$ be a probability space. Let $X_{1}^{(n)},\ldots,X_{d}^{(n)}$ be a sequence of $d$-tuples of independent random matrices of size $n\times n$ with independent and identically distributed Gaussian entries over $\Omega$ with mean zero and variance $\frac{1}{n}$. Then, given any $f\in\mathscr{M}_{d}$, if $f(s_{1},\ldots,s_{d})$ is defined, then $f(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))$ is well defined for sufficiently large $n\in\mathbb{N}$ and almost every $\omega\in\Omega$, and
@@ 2.1 Non-commutative functions
One can topologize $\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ in several ways. The disjoint union topology was first considered by Taylor [47], the uniform topology was introduced in [25], and the free topology was studied in [2]. It is a surprising result [1, 25] that for an NC function defined on a disjoint-union open set, local boundedness automatically implies the function is holomorphic. Moreover, any such locally bounded NC function is analytic in the sense that it admits a certain Taylor-type power series expansion, called a Taylor–Taylor series about every point in its NC domain with non-zero radius of convergence. In particular, the Taylor–Taylor series of any such function at the origin, $0=(0,\cdots,0)\in\mathbb{C}^{1\times d}$ of the NC universe is a free formal power series with complex coefficients [25, Corollary 4.4].
@@
Taking $Y=0$ at level one of the NC universe yields the NC unit row-ball. The uniform topology on $\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ is then the topology generated by the sub-base of open sets consisting of all uniform NC unit row-balls centered at any point $Y\in\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ [25, Section 7.2]. Again, in [25, Section 7.2], the uniform topology can be defined with respect to any operator space structure on $\mathbb{C}^{d}$; we have chosen to work with the row operator space structure. In this paper, we study algebras of uniformly bounded NC functions defined on uniformly open NC row-balls centered at 0. Again, since such functions are bounded, hence locally bounded in the uniform topology, they are holomorphic in the sense that they are Fréchet differentiable about any point in their uniformly open domains, and they are analytic in the sense that they have a Taylor–Taylor series power series expansion about any point in their domains with non-zero radius of convergence [25, Chapter 7].
In [41, Section 3], Popescu introduced the algebra of all complex free formal power series with radius of convergence at least $R\in(0,+\infty]$:
$\mathscr{O}_{d}(R):=\left\{\left.f\in\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle\right|\ R_{f}\geq R\right\},$ (2.1)
@@
It is not hard to check that each such power series defines an NC function on $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ that is bounded on every $r\cdot\overline{\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}}$ for $0<r<R$. Conversely, if we have an NC function $f$ defined on $\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ that is bounded on all row sub-balls of finite radii, then by [41, Theorem 1.1 and Section 3], or by a special case of [25, Corollary 7.26 and Theorem 8.11], the Taylor–Taylor expansion of $f$ around the origin is a free formal power series that converges absolutely and uniformly on each such ball, and hence defines a uniformly analytic NC function in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. Therefore, by [41, Theorem 1.1], $f\in\mathscr{O}_{d}(R)$. In particular, the ring of uniformly entire NC functions, $\mathscr{O}_{d}=\mathscr{O}_{d}(\infty)$, can be defined as the sub-ring of $\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ consisting of free FPS with infinite radius of convergence, or equivalently, as the algebra of NC functions defined on the entire NC universe which are bounded on any NC row-ball of finite radius. It is immediate from the definition that every NC function in $\mathscr{O}_{d}$ can be evaluated on any $d$-tuple of linear operators acting on a complex, separable Hilbert space simply by substituting the operators for the formal variables in the free FPS expansion.
@@
J.L. Taylor developed a theory of “localizations” of the free algebra $\mathbb{C}\langle\mathfrak{z}\rangle$ in [47]. Taylor’s localization is a topological algebra over the free algebra with a property that is today known as a homological ring epimorphism [17, Definition 4.5]. (This is a different notion of “localization” than we consider in this paper; our universal localizations of semifirs are universal skew fields of fractions.) One family of such homological ring epimorphisms that Taylor proposes is the family of algebras $S(r)$ for $r=(r_{1},\cdots,r_{d})$ a $d$-tuple of positive numbers. Here,
@@
We write for a word $\alpha$ in the alphabet $\{1,\ldots,d\}$, $\mathfrak{z}^{\alpha}=\mathfrak{z}_{\alpha_{1}}\cdots\mathfrak{z}_{\alpha_{n}}$ with the obvious evaluation on numbers and matrices. In this paper, we will show that in many regards the different mode of convergence introduced by Popescu [41] is more amenable to analysis than the one proposed by Taylor (see also [29] and [25]).

### v2

@@ Abstract
We show, in particular, that the ring of (uniformly) entire non-commutative (NC) functions in $d\in\mathbb{N}$ non-commuting variables, $\mathscr{O}_{d}$, is a semifir. Every finitely–generated right (or left) ideal in $\mathscr{O}_{d}$ is closed, which yields an analytic extension of G. M. Bergman’s nullstellensatz for the free algebra. Any semifir admits a universal skew field of fractions; applying this to $\mathscr{O}_{d}$ yields the universal skew field of “NC meromorphic expressions”, $\mathscr{M}_{d}$. We show that any $f\in\mathscr{M}_{d}$ has a well-defined domain and evaluations in a large class of stably-finite topological algebras, including finite $C^{*}$-algebras, extending a result of Cohn for NC rational functions. As an application, we extend an “almost sure convergence” result of Haagerup and Thorbjørnsen for free polynomials evaluated on tuples of random matrices to the setting of NC meromorphic expressions.
@@ 1 Introduction
Non-commutative function theory originated in the work of J.L. Taylor in multivariate spectral theory and functional calculus for tuples of non-commuting operators [50, 49]. A non-commutative function in several non-commuting variables is a function on tuples of square matrices of any fixed, finite size, taking values in square matrices, and which has natural invariance properties enjoyed by free polynomials, $p\in\mathbb{C}\langle\mathfrak{z}\rangle=\mathbb{C}\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle$. Here, $\mathbb{C}\langle\mathfrak{z}\rangle$ denotes the free algebra over $\mathbb{C}$ on $d$ generators. We will call elements of $\mathbb{C}\langle\mathfrak{z}\rangle$ non-commutative or free polynomials. A function, $f$, on a direct sum-closed subset of the complex NC universe of $d$-tuples of square matrices over $\mathbb{C}$ of any fixed finite size is an NC function if it: (i) respects the grading (matrix size), (ii) respects direct sums and (iii) respects joint similarities, see Subsection 2.1 for definitions and details. It is readily checked that any $p\in\mathbb{C}\langle\mathfrak{z}\rangle$ is an NC function on the entire NC universe. These three axioms defining NC functions are surprisingly rigid: Any NC function that is ‘continuous’ or even ‘locally bounded’ in an ‘open’ NC set is automatically holomorphic in the sense that it is Fréchet differentiable at any point in its domain and analytic in the sense that it has a convergent Taylor-type power series expansion about any point of its domain with non zero “radius of convergence”. (In order to precisely define the punctuated terms in the previous sentence, we need to introduce a topology on the NC universe, which we will do in Subsection 2.1. The topology that is most relevant for our investigations here is called the uniform topology [26, Section 7.2].)
@@
where $\omega=i_{1}\cdots i_{n}$ is any word comprised of letters $i_{j}$, chosen from the alphabet, $i_{j}\in\{1,\cdots,d\}$. Here, $\mathbb{W}_{d}$ denotes the free monoid, the set of all such words in $d$ letters. We also consider the empty word, $\emptyset$, consisting of no letters, and the free monomials, $\mathfrak{z}^{\omega}$, are then defined in the obvious way: $\mathfrak{z}^{\omega}=\mathfrak{z}_{i_{1}}\cdots\mathfrak{z}_{i_{n}}$ if $\omega=i_{1}\cdots i_{n}$ and $\mathfrak{z}^{\emptyset}=:1$. G. F. Popescu [43] introduced an analogue of the Cauchy–Hadamard radius of convergence formula for such free FPS, $f$, given by
@@
see also [30] and [26]. In the above, $|\omega|\in\mathbb{N}\cup\{0\}$ denotes the length of $\omega\in\mathbb{W}_{d}$, i.e., the number of letters it contains. Given any $R\in(0,+\infty]$, we study the rings, $\mathscr{O}_{d}(R)$, of all free formal power series with radius of convergence at least $R$. We write $\mathscr{O}_{d}:=\mathscr{O}_{d}(+\infty)=\cap_{R>0}\mathscr{O}_{d}(R)$, for the ring of NC entire functions. As proven in [43, Theorem 1.1], if $f\in\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$, has radius of convergence $R_{f}>0$, then $f(X)$ converges absolutely and uniformly in an NC row-ball, $R_{f}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, centred at the origin, $0=(0,\cdots,0)\in\mathbb{C}^{1\times d}$, of the $d$-dimensional NC universe, consisting of all square matrix $d$-tuples, $X=(X_{1},\cdots,X_{d})\in\mathbb{C}^{n\times n}\otimes\mathbb{C}^{1\times d}$, which define bounded linear maps from $\mathbb{C}^{n}\otimes\mathbb{C}^{d}$ into $\mathbb{C}^{n}$, with norm less than $R_{f}$. Hence, any $f\in\mathscr{O}_{d}(R)$ defines a (uniformly) analytic NC function in $R_{f}\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. In particular, $\mathscr{O}_{d}$ consists of holomorphic and analytic non-commutative functions on the entire NC universe and is, in a very natural sense, a canonical multivariate extension of the ring of entire functions in $\mathbb{C}$ to several non-commuting variables.
In pure non-commutative algebra, semi-free ideal rings were introduced by P.M. Cohn in his study of localizations and universal skew fields of fractions of non-commutative rings, see [10] for a full development of this theory. A ring, $\mathscr{R}$, is a semi-free ideal ring (semifir for short), if every finitely generated right ideal of $\mathscr{R}$ is free as a right $\mathscr{R}$-module. This notion is left-right symmetric. Cohn showed that every semi-free ideal ring admits a (necessarily unique) universal skew field of fractions. There are not many known examples of semi-free ideal rings, and many non-commutative rings do not have universal localizations. (Certainly, any ring with zero divisors cannot be embedded into a skew field.) Among the few known examples are the free algebra, the ring $\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle=\mathbb{C}\langle\!\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle\!\rangle$ of formal power series in several non-commuting variables, the sub-ring of algebraic free FPS, the group ring of the free group and group rings over locally free groups (see, for example, [10] and [14]). Some more recent examples of universal localizations and of semifirs were constructed by Klep, Vinnikov, and Volčič in [27, 28]. The first main result of this paper provides a new family of examples of semifirs of uniformly analytic NC functions:
@@
It follows that each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$ has a universal skew field of fractions, which we denote by $\mathscr{M}_{d}(R)$. In particular, we will call $\mathscr{M}_{d}:=\mathscr{M}_{d}(\infty)$ the field of (global) NC (uniformly) meromorphic expressions – c.f. the local field of (uniformly) meromorphic NC germs at $0\in\mathbb{C}^{1\times d}$ investigated in [27]. Our proof of the above theorem relies heavily on the operator–algebraic “Non-commutative Hardy space theory” initiated by Popescu and Davidson–Pitts, in which the full Fock space over $\mathbb{C}^{d}$ is interpreted as the “free Hardy space” of square–summable formal power series in $d$ NC variables [43, 42, 39, 13]. In particular, we will employ the non-commutative inner–outer factorization [13, 35, 37] for the free Hardy space, and extensions of methods in [25] in conjunction with the structure of $\mathscr{O}_{d}(R)$ as an inverse limit of non-commutative disk algebras on balls of increasing radii [5].
We further obtain an analytic extension of a non-commutative nullstellensatz for the free algebra due to G. M. Bergman to each of our semifirs, $\mathscr{O}_{d}(R)$, see [24]. To this end, define the directional variety of $S\subseteq\mathscr{O}_{d}(R)$ as
@@
In Section 6, we extend a theorem of P.M. Cohn for evaluation of non-commutative rational functions in any stably-finite algebra. Here, the free skew field of all NC rational functions, $\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,:=\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$, is the universal skew field of fractions of the free algebra, $\mathbb{C}\langle\mathfrak{z}\rangle=\mathbb{C}\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle$ [4, 10]. Namely, we prove that NC meromorphic elements in $\mathscr{M}_{d}$ have well-defined evaluations and domains in any stably finite algebra, $\mathscr{A}$, that admits well-defined unital ring homomorphisms $\varphi:\mathscr{A}\rightarrow\mathscr{O}_{d}$; in particular in any stably finite $C^{*}$-algebra. This result enables us to develop concrete applications of our results to D.-V. Voiculescu’s free probability theory and random matrix theory. Namely, we extend the results of Haagerup and Thorbjørnsen for the free algebra [21] and of Yin [56] for the free skew field, to the larger field of NC meromorphic expressions. To describe these results in more detail, let $s_{1},\cdots,s_{d}$ denote the real parts of the left creation operators on the full Fock space over $\mathbb{C}^{d}$. This $d$-tuple is the prototypical example of free semicircular random variables in free probability theory.
@@
Let $(\Omega,P)$ be a probability space. Let $X_{1}^{(n)},\cdots,X_{d}^{(n)}$ be a sequence of $d$-tuples of independent random matrices of size $n\times n$ with independent and identically distributed Gaussian entries over $\Omega$ with mean zero and variance $\frac{1}{n}$. Then, given any $f\in\mathscr{M}_{d}$, if $f(s_{1},\cdots,s_{d})$ is defined, then $f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$ is well defined for sufficiently large $n\in\mathbb{N}$ and almost every $\omega\in\Omega$, and
@@ 2.1 Non-commutative functions
One can topologize $\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ in several ways. The disjoint union topology was first considered by Taylor [49], the uniform topology was introduced in [26], and the free topology was studied in [2]. It is a surprising result [1, 26] that for an NC function defined on a disjoint-union open set, local boundedness automatically implies the function is holomorphic. Moreover, any such locally bounded NC function is analytic in the sense that it admits a certain Taylor-type power series expansion, called a Taylor–Taylor series about every point in its NC domain with non-zero radius of convergence. In particular, the Taylor–Taylor series of any such function at the origin, $0=(0,\cdots,0)\in\mathbb{C}^{1\times d}$ of the NC universe is a free formal power series with complex coefficients [26, Corollary 4.4].
@@
Taking $Y=0$ at level one of the NC universe yields the NC unit row-ball. The uniform topology on $\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ is then the topology generated by the sub-base of open sets consisting of all uniform NC unit row-balls centered at any point $Y\in\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ [26, Section 7.2]. Again, in [26, Section 7.2], the uniform topology can be defined with respect to any operator space structure on $\mathbb{C}^{d}$; we have chosen to work with the row operator space structure. In this paper, we study algebras of uniformly bounded NC functions defined on uniformly open NC row-balls centered at 0. Again, since such functions are bounded, hence locally bounded in the uniform topology, they are holomorphic in the sense that they are Fréchet differentiable about any point in their uniformly open domains, and they are analytic in the sense that they have a so-called Taylor–Taylor series power series expansion about any point in their domains with non-zero radius of convergence [26, Chapter 7].
In [43, Section 3], Popescu introduced the algebra of all complex free formal power series with radius of convergence at least $R\in(0,+\infty]$:
$\mathscr{O}_{d}(R):=\left\{\left.f\in\mathbb{C}\langle\!\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle\!\rangle\right|\ R_{f}\geq R\right\},$ (2.1)
@@
It is not hard to check that each such power series defines an NC function on $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ that is bounded on every $r\cdot\overline{\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}}$ for $0<r<R$. Conversely, if we have an NC function $f$ defined on $\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ that is bounded on all row sub-balls of finite radii, then by [43, Theorem 1.1 and Section 3], or by a special case of [26, Corollary 7.26 and Theorem 8.11], the Taylor–Taylor expansion of $f$ around the origin is a free formal power series that converges absolutely and uniformly on each such ball, and hence defines a uniformly analytic NC function in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. Therefore, by [43, Theorem 1.1], $f\in\mathscr{O}_{d}(R)$. In particular, the ring of uniformly entire NC functions, $\mathscr{O}_{d}=\mathscr{O}_{d}(\infty)$, can be defined as the sub-ring of $\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ consisting of free FPS with infinite radius of convergence, or equivalently, as the algebra of NC functions defined on the entire NC universe which are bounded on any NC row-ball of finite radius. It is immediate from the definition that every NC function in $\mathscr{O}_{d}$ can be evaluated on any $d$-tuple of linear operators acting on a complex, separable Hilbert space simply by substituting the operators in place of the formal variables in the free FPS expansion.
@@
J.L. Taylor developed a theory of “localizations” of the free algebra $\mathbb{C}\langle\mathfrak{z}\rangle$ in [49]. Taylor’s localization is a topological algebra over the free algebra with a property that is today known as a homological ring epimorphism [18, Definition 4.5]. (This is a different notion of “localization” than we consider in this paper; our universal localizations of semifirs are universal skew fields of fractions.) One family of such homological ring epimorphisms that Taylor proposes is the family of algebras $S(r)$ for $r=(r_{1},\cdots,r_{d})$ a $d$-tuple of positive numbers. Here,
@@
In this paper, we will show that in many regards the different mode of convergence introduced by Popescu [43] is more amenable to analysis than the one proposed by Taylor (see also [30] and [26]).